# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v1

@@ 6.5 Other Pre-trained Weights
To more comprehensively explore the impact of $f_{0}^{*}$ in Eq. 3 on the final performance of PEFT-CL, extensive experiments using other pre-trained weights for ViT-B/16 are conducted. To ensure absolute fairness, the hyper-parameters and training strategies involved during their training are kept completely consistent, with only the backbone parameters differing. The results, as shown in Table VII, reveal several key insights.
Firstly, self-supervised methods exhibit significant variability, with iBOT ImageNet-22K [101] achieving the highest incremental accuracy on CIFAR-100 and ImageNet-A. This indicates a positive correlation between sample size, model generalizability, and resistance to forgetting, as articulated in lemma 3. In addition, the autoregressive self-supervised method MAE exhibits extreme disadvantages in both performance on individual tasks and resistance to forgetting. The reasons for this are not yet known, leaving room for further exploration.
Secondly, traditional supervised pre-trained weights, consistently demonstrate superior performance across all datasets, clearly outperforming other customized supervisory weights. This trend suggests that excessive pre-training strategies may not necessarily enhance model generalizability in PEFT-CL, despite their potential advantages under standard training paradigms. Furthermore, CLIP-Vision, which relies solely on visual information, only achieves state-of-the-art results on ImageNet-R. This outcome does not negate the value of more extensive sample sets. Instead, it implies that the dual-modal (Vision+Text) pre-training paradigm may not fully capitalize on the benefits of extensive training samples when employed solely within the visual modality.
@@ 7 Conclusion
In this study, we adopt an NTK perspective to analyze PEFT-CL tasks, elucidating model behavior and generalization gaps in sequential task learning. Our analysis identifies crucial factors affecting PEFT-CL effectiveness, particularly through the dynamics of task interplay and task-specific generalization gaps. We recommend strategies to mitigate these gaps, such as expanding sample sizes, enforcing task-level feature constraints, and refining regularization techniques. These strategies inform architectural and optimization adjustments aimed at enhancing the generalization capabilities and efficiency of PEFT-CL models across diverse tasks. This research significantly advances the theoretical framework and provides practical guidelines for improving PEFT-CL methodologies, contributing to both academic insights and real-world applications.
Appendix A NTK Dynamics in PEFT-CL
Initially, we concentrate on analyzing the least squares loss associated with the optimization of consecutive tasks $\tau$ and $\tau-1$. This involves quantifying the classification loss attributable to variations in the sub-network components’ parameters, which is expressed as follows:
@@ Appendix A NTK Dynamics in PEFT-CL
Here, $D_{\tau}$ refers to the data subset associated with the $\tau$-th task, where $X$ and $Y$ are the input images and corresponding labels, respectively. The term $\phi_{\tau}(\cdot)$ denotes the Jacobian matrix relevant to task $\tau$ for the inputs $X$. At the onset of a task’s optimization, the sub-network component parameters inherit parameters from the preceding task, setting the initial states for optimizing $f_{\tau}(\cdot)$ and $p_{\tau}$ as $f_{\tau-1}^{*}(\cdot)$ and $p_{\tau-1}^{*}$, respectively.
@@
Appendix B Task-Interplay Generalization in PEFT-CL
@@
In accordance with Mercer’s Theorem [55], within an ideal RKHS, the NTK can be expressed as an infinite sum of orthogonal basis functions and eigenvalues:
@@
Appendix C Task-Intrinsic Generalization in PEFT-CL
@@
In an approach analogous to the solution process for NTK Dynamics discussed in Appendix A, we construct a kernel regression error for the weight matrix $w$:
@@
Drawing from [72] and [10], we aim to determine the dynamic changes of $U_{\rho,\gamma}$. Introducing auxiliary variable $z$ and data quantity variable $s$, $U_{\rho,\gamma}$ can be represented as:
@@
At this stage of the analysis, by applying the Woodbury Matrix Inversion Formula [70], we derive the following expression:
@@
Appendix D Task Segmentation
In Tables VIII and IX, we outline the class order for CIFAR100, ImageNet-R, and ImageNet-A for each seed configuration. All subsequent task segmentations adhere to these class orders. The method to establish this class order involves setting the random seed and executing a random permutation of the class indices during task segmentation definition. The following code snippet illustrates this process:
@@ Appendix D Task Segmentation
TABLE VIII: The class order for each seed on CIFAR100 determines all subsequent task segmentations.
@@
TABLE IX: The class order for each seed on ImageNet-A and ImageNet-R determines all subsequent task segmentations.
@@
Appendix E Platonic Representation in PEFT-CL

### v2

@@ 6.6 Other Pre-trained Weights
To more comprehensively explore the impact of $f_{0}^{*}$ in Eq. 3 on the final performance of NTK-CL, extensive experiments using other pre-trained weights for ViT-B/16 are conducted. To ensure absolute fairness, the hyper-parameters and training strategies involved during their training are kept completely consistent, with only the backbone parameters differing. The results, as shown in Table XII, reveal several key insights.
Firstly, self-supervised methods exhibit notable variability in performance across various continual tasks. Among them, iBOT ImageNet-22K [110] achieves the highest incremental accuracy on both CIFAR-100 and ImageNet-A, indicating a positive correlation between the scale of pre-training data, model generalization, and resistance to forgetting, as discussed in lemma 3. In contrast, the masked modeling generative method MAE demonstrates significant limitations, with inferior performance in both task-specific accuracy and knowledge retention. This deficiency is primarily attributed to its pixel-level masked reconstruction objective, which emphasizes low-level structural recovery at the expense of learning semantically discriminative features. As a result, MAE fails to maintain sufficient class separability, reducing its effectiveness in NTK-CL. Secondly, supervised pre-trained weights consistently deliver superior performance across all evaluated datasets, significantly outperforming both self-supervised and customized supervised alternatives. This suggests that excessive specialization in pre-training objectives does not necessarily enhance generalization in PEFT-CL scenarios. Finally, CLIP-Vision, despite relying solely on visual modality input, achieves state-of-the-art performance exclusively on the ImageNet-R dataset, while exhibiting slight limitations on other benchmarks. We attribute this phenomenon to its alignment with the semantic complexity inherent in ImageNet-R. To further investigate the causes of performance variation, particularly MAE and CLIP-Vision, we provide detailed visualization analyses in Appendix I.
@@ 7 Conclusion
In this study, we adopt an NTK perspective to analyze PEFT-CL tasks, elucidating model behavior and generalization gaps in sequential task learning. Our analysis identifies crucial factors affecting PEFT-CL effectiveness, particularly through the dynamics of task-interplay and task-specific generalization gaps. We recommend strategies to mitigate these gaps, such as expanding sample sizes, enforcing task-level feature constraints, and refining regularization techniques. These strategies inform architectural and optimization adjustments, enhancing model generalization while advancing their theoretical and practical foundations.
8 Future Work and Discussions
With the emergence of pre-trained Large Language Models (LLMs), a fundamental challenge is extending the NTK-CL framework to encompass both LLMs and Multimodal Large Language Models (MLLMs/Omni-Models). Although several preliminary approaches have been proposed [14, 108, 67, 71, 84, 100], they predominantly focus on simplified architectures, such as T5, and have yet to demonstrate scalability or efficacy on more sophisticated LLMs and generalist Omni-Models. A detailed discussion of these limitations is provided in Appendix M. Additionally, although generative self-supervised pre-training schemes (e.g., MAE) achieve strong generalization across some other domains, their deployment in PEFT-CL settings exposes limitations, particularly the issue of semantic indistinguishability, which remains an open problem and warrants further investigation. Lastly, future work should prioritize developing theoretically grounded Bayesian hyper-parameter search algorithms that, like Dynamic Loss Scaling, introduce minimal overhead while ensuring rigorous mathematical guarantees.
Appendix A Prelude to the Study
In the realm of PEFT-CL, we initiate with the foundational model $f_{0}^{*}$, where the superscript signifies parameters optimized to their prime configuration, distinguishing them from those still under optimization. Our objective is to adeptly modify the feature space from $f_{0}^{*}$ for each specific task $i$ (represented as $f_{i}^{*}$), by finely adjusting the optimizable subnetwork parameters $p_{i}$. This critical adaptation ensures that alterations in shared subnetwork components across different tasks do not lead to excessive catastrophic forgetting, thereby safeguarding the model’s generalizability.
To methodically investigate PEFT-CL, we conceptualize a sequence of $T$ tasks, each optimizing subnetwork parameters $p_{\tau}^{*}$ for $1\leq\tau\leq T$. This strategic adjustment of $p_{\tau}^{*}$ enables the model to achieve an optimal state $f_{\tau}^{*}$, thereby generating a specialized feature space for each task.
Expanding beyond traditional heuristic approaches [88, 87, 76, 39, 109] prevalent in PEFT-CL for adjusting subnetwork components, our analysis delves into the training dynamics and explores the potential for reducing generalization gaps through the lens of NTK. This rigorous analysis helps pinpoint necessary adjustments to minimize generalization gaps and maximize model performance across diverse tasks. Grounded in seminal theories and contemporary studies in generalization dynamics [20, 11, 9, 40, 4], we introduce advanced tools for assessing task interplay and specific generalization gaps, as detailed in theorem 1 and theorem 2.
For comprehensive clarity in theoretical discourse, we segment our discussion into three distinct parts: analyzing NTK dynamics specific to PEFT-CL (Appendix B), evaluating inter-task generalization gap (Appendix C), and scrutinizing intra-task generalization gap (Appendix D). These segments collectively aim to provide a deep understanding of the underlying mechanisms influencing PEFT-CL performance, thereby informing better implementation practices in this field.
Appendix B NTK Dynamics in PEFT-CL
Initially, we concentrate on analyzing the least squares loss associated with the optimization of consecutive tasks $\tau$ and $\tau-1$. This involves quantifying the classification loss attributable to variations in the subnetwork components’ parameters, which is expressed as follows:
@@ Appendix B NTK Dynamics in PEFT-CL
Here, $D_{\tau}$ refers to the data subset associated with the $\tau$-th task, where $X$ and $Y$ are the input images and corresponding labels, respectively. The term $\phi_{\tau}(\cdot)$ denotes the Jacobian matrix relevant to task $\tau$ for the inputs $X$. At the onset of a task’s optimization, the subnetwork component parameters inherit parameters from the preceding task, setting the initial states for optimizing $f_{\tau}(\cdot)$ and $p_{\tau}$ as $f_{\tau-1}^{*}(\cdot)$ and $p_{\tau-1}^{*}$, respectively.
@@
Appendix C Task-Interplay Generalization in PEFT-CL
@@
In accordance with Mercer’s Theorem, within an ideal RKHS, the NTK can be expressed as an infinite sum of orthogonal basis functions and eigenvalues:
@@
Appendix D Task-Intrinsic Generalization in PEFT-CL
@@
In an approach analogous to the solution process for NTK Dynamics discussed in Appendix B, we construct a kernel regression error for the weight matrix $w$:
@@
Drawing from [11], we aim to determine the dynamic changes of $U_{\rho,\gamma}$. Introducing auxiliary variable $z$ and data quantity variable $s$, $U_{\rho,\gamma}$ can be represented as:
@@
At this stage of the analysis, by applying the Woodbury Matrix Inversion Formula, we derive the following expression:
@@
Appendix E Datasets and Experimental Configurations
Datasets: Specifically, we utilize the CIFAR-100 dataset [42], which consists of 60,000 32x32 color images distributed across 100 classes. To align with the input requirements of the pre-trained ViT model, the images are resized to 224x224 pixels and organized into 10 tasks, each comprising 10 classes. Additionally, the ImageNet-R dataset [87] is employed, which extends the original ImageNet by incorporating artistic renditions, cartoons, and stylized interpretations for 200 classes, structured into 10 tasks with 20 classes each, featuring 24,000 training and 6,000 test images. The ImageNet-A dataset [30] further evaluates the generalization of models against adversarial and out-of-distribution samples, consisting of 7,500 images from 200 classes, partitioned into 10 tasks. The DomainNet dataset [63], a large-scale domain adaptation resource, is also utilized. It comprises six distinct domains—Clipart, Infograph, Painting, Quickdraw, Real, and Sketch—totaling 423,506 images across 345 categories. These are organized into 15 tasks, each containing 23 classes, to thoroughly test cross-domain generalization. Unlike prior studies, such as DAP [39], which focuses on the Real domain, and CODA-Prompt [76], which examines a limited five-task sequence within the Real domain, our study encompasses all six domains in a structured 15-task sequence. This approach establishes a more comprehensive benchmark for the continual domain adaptation.
Furthermore, we incorporate additional datasets, including Oxford Pets [61], EuroSAT [29], PlantVillage [34], VTAB [103], and Kvasir [65], as detailed in Table I. This extensive dataset selection underscores the robustness, generalization, and adaptability of our framework across a wide range of visual recognition tasks, thereby validating its efficacy in addressing domain-specific challenges.
Evaluation Metrics: Following the established benchmark protocol in [72], we evaluate the model’s effectiveness using $A_{\tau}$, which signifies the accuracy post the $\tau$-th training stage. Notably, we employ $A_{T}$—the performance metric at the termination of the final stage—and $\bar{A}=\frac{1}{T}\sum_{\tau=1}^{T}A_{\tau}$, which calculates the average accuracy over all incremental stages. These metrics are selected as the principal measures of model performance, providing a holistic view of its efficacy and stability throughout the training process.
Appendix F Task Segmentation
In Tables XIII and XIV, we outline the class order for CIFAR100, ImageNet-R, and ImageNet-A for each seed configuration. All subsequent task segmentations adhere to these class orders. The method to establish this class order involves setting the random seed and executing a random permutation of the class indices during task segmentation definition. The following code snippet illustrates this process:
@@ Appendix F Task Segmentation
TABLE XIII: The class order for each seed on CIFAR100 determines all subsequent task segmentations.
@@
TABLE XIV: The class order for each seed on ImageNet-A and ImageNet-R determines all subsequent task segmentations.
@@
Appendix G Platonic Representation in PEFT-CL